there are two urns A and B, that contain black and white marbles. Urn a contains 3 white marbles and 2 black marbles. Urn b contains 24 white marbles and 16 black marbles. If your reach into one of the two urns at random and pull out a white marble, what is the probability that your reached into urn A. Please express your answer to nearest whole percent
Answer #1

Please find below the answer. Don't hesitate to give a "thumbs up" for the answer in case you're satisfied with the answer.

So, the probability that it is a white ball = (1/2)*(3/5)+ (1/2)*(24/40)

The probability that the white ball came from urn A i s:

= (1/2)*(3/5) / (1/2)*(3/5)+ (1/2)*(24/40)

= (.6)/ (.6+.6)

= 1/2

= .50

This is basically 50%